Circular No. 8989

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1721 AQUILAE = NOVA AQUILAE 2008 H. Yamaoka, Kyushu University, reports the discovery by K. Itagaki

(Yamagata, Japan) of a variable star on his CCD survey image taken on Sept. 22.5 UT with a 0.21-m reflector; a confirming unfiltered CCD image taken on Sept. 22.586 with a 0.60-m reflector yields mag ∼ 14.0 and the following precise position: α = 19^{h}06^{m}28^{s}.58, δ = +7^{o}06′44′′.3 (equinox 2000.0). Additional magnitudes from Itagaki (via S. Nakano): 2007 Nov. 2.396, [17.0; Sept. 22.586, 14.0; 24.489, 14.1; 27.445, 15.8; 29.529, 16.5; Oct. 1.38759, 16.7; 2.421, 16.9; 3.433, 17.1. The 2MASS catalogue contains a very faint star with position end figures 28^{s}.60, 44′′.5. Following posting on the Central Bureau’s unconfirmed-objects webpage, V. Nevski (Vitebsk, Belarus, 0.30-m reflector) reports that an unfiltered CCD image obtained on Sept. 22.8 yields red mag 14.0 and position end figures 28^{s}.59, 44′′.6 for the variable, adding that nothing is visible at this position on Digitized Sky Survey images from 1951 July 5 and 1987 July 30 (limiting mag ≈ 20). E. Y. Hsiao, M. L. Graham, C. J. Pritchet, and D. D. Balam, University of Victoria, report that a spectrogram (range 390–703 nm, resolution 0.3 nm), obtained on Sept. 23.16 using the 1.82-m Plaskett Telescope, exhibits a strongly increasing continuum from 600 to 703 nm with a broad and flat-topped emission (HWZI ∼ 2700 km/s, 6.5 nm) at the position of Hα; additional details are given on CBET 1512. E. Kazarovets reports that the GCVS team assigns the designation V1721 Aql to this object.

L. A. Helton and C. E. Woodward, University of Minnesota; and K. Vanlandingham and G. J. Schwarz, West Chester University, report on Boller-and-Chivens spectroscopic observations (range 380–900 nm; resolution ∼ 0.28 nm) of this nova that were obtained at the Steward Observatory Bok 2.29-m telescope on Sept. 25.19 and 25.25 UT. Hα exhibits a very broad (FWHM = 6450 km/s) tri-peaked profile with emission components at +3020, −140, and −3550 km/s. Permitted oxygen makes an appearance at 777.3 and 844.6 nm with broad (FWHM ∼ 7600 and 7500 km/s, respectively), tri-peaked structure similar to Hα. The spectrum is heavily reddened with a complete absence of any emission features blueward of Hα. The high reddening is confirmed by the extinction maps of Schlegel et al. (1998, Ap.J. 500, 525), which indicate exceedingly high reddening along this line-of-sight but suffer from very large uncertainties since b ∼ 0. Comparison with other novae at a similar early evolutionary state imply E(B–V) = 3 and a likely distance of 5 kpc, assuming a maximum absolute V magnitude of −9.

2008 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green